OpenStudy (anonymous):

Verify the identity. cos(pi/2-Beta)=sinBeta

OpenStudy (anonymous):

\[\cos a-b=\cos a \cos b+\sin a \sin b\] puting here a as 90 and b as beta u will get the result

OpenStudy (anonymous):